# Variational quantum state preparation for quantum-enhanced metrology in noisy systems

### v1

@@ III.1.3 Additional diagnostic quantities
In addition to the fidelity to the GHZ state and the squeezing parameter, it is insightful to consider additional quantities. First, we compute the collectivity of the prepared states [see Fig. 3], $\langle\hat{S}^{2}\rangle$ which should be compared against the maximal value of $S(S+1)$ with $S=N/2$ for a collective state. Both OAT and TAT unitary dynamics preserve the collectivity as they feature infinite-range interactions, so that any deviation from the maximum value of $\langle\hat{S}^{2}\rangle$ should be attributable to the single-qubit dephasing noise. The collectivity is most useful examined alongside a second quantity, the classical Fisher information (CFI), which can be obtained as,
$F_{c}(\theta)=\lim_{\theta\to 0}\frac{4}{\theta^{2}}\sum_{m_{z}}\left(\sqrt{P(m_{z};\theta)}-\sqrt{P(m_{z};\theta=0)}\right)^{2}$ (10)
@@
where $P(m_{z})$ is the distribution function for the collective spin projection $m_{z}$ along $\hat{z}$, and $F_{c}$ is bounded from above by the QFI, $F_{c}\leq F_{Q}$. In the case of unitary OAT and TAT dynamics, the CFI always saturates the QFI when optimized over all possible collective measurements. However, we observe $F_{c}$ is less than the QFI across parts of both the perturbative and squeezed-like regimes, particularly when the collectivity strays appreciably from $S(S+1)$. This suggests that more sophisticated measurements, such as local (site-resolved) spin observables are necessary to exhaust the metrological potential of the states we preserve, and in particularly highlights that states in the “squeezed” regime – a name that is motivated by a simple analysis of collective spin fluctuations – require more nuanced treatment in general.
III.2 Ising, XX, f-TAT
@@
Figure 4: Diagnostic quantities calculated for optimally prepared states as functions of dephasing strength $\gamma$ for (a) Ising, (b) XX and (c) f-TAT models. All data is for $N=8$ and $\alpha=6$. (Top subplots) Normalized classical Fisher information ($F_{c}/N^{2}$, squares), inverse Wineland squeezing parameter $(N\xi^{2}_{s})^{-1}$, tri-prongs) and the normalized QFI ($F_{Q}/N^{2}$, circles, diamonds and triangles, respectively). (Lower subplots) The overlap of the optimal state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ with a GHZ state ($\mathcal{F}_{\mathrm{GHZ}}$, diamonds) and the normalized collectivity ($\langle\hat{S}^{2}\rangle/S(S+1)$, crosses). Vertical lines in all panels indicate $\gamma^{(1)}$ [absent for (a)] and $\gamma^{(2)}$. Below each set of panels we show Husimi phase space distributions for representative quantum states in each regime (see Appendix A for details of distribution). Bright regions indicate larger probability density. For the GHZ-like states we show two perspectives of the same quantum state, rotated by $\pi/2$ about indicated axes.
@@
Equipped with an understanding of the results for the infinite-ranged OAT and TAT entangling Hamiltonians, we now generalize to the finite-ranged models given in Eqs. (2)-(4). For simplicity, we focus on the specific coupling exponent $\alpha=6$ in the following, and will present additional results investigating the role of interaction range in a more detailed follow-up work Zuñiga Castro et al. (ming).
@@
In Fig. 2(b), we plot the maximal QFI and optimal parameters $\mathbf{x}_{\mathrm{opt}}$ as a function of dephasing strength $\gamma$ with $N=8$ qubits. Overall, we observe a similar delineation of the QFI and $\mathbf{x}_{\mathrm{opt}}$ into different regimes set by the dephasing strength, with some important caveats. Firstly, for the XX and f-TAT models, we define an analogous perturbative regime for $\gamma<\gamma^{(1)}$ (see vertical lines in panels), where the transition $\gamma^{(1)}$ is again defined by an abrupt drop in the interaction time characterized by $\theta_{I}$ and similarly sharp changes in the rotation angles/axes. Similarly, we define a squeezed-like regime for $\gamma^{(1)}<\gamma<\gamma^{(2)}$, with the upper bound on the dephasing strength $\gamma^{(2)}$ set via the uncorrelated regime that emerges when $\theta_{I}$ finally reaches zero. On the other hand, the Ising model presents distinct behaviour. In particular, we observe an absence of the perturbative regime, as the VQC parameters change smoothly throughout a large squeezed-like regime $\gamma<\gamma^{(2)}$, until $\theta_{I}$ vanishes at the boundary of the uncorrelated regime that exists for $\gamma\geq\gamma^{(2)}$.
@@
While our preliminary identification of the perturbative, squeezed-like and uncorrelated regimes is supported by an analysis of the optimal parameters $\mathbf{x}_{\mathrm{opt}}$, which show some qualitative similarities to OAT (Ising and XX) and TAT (f-TAT), we provide stronger evidence through quantities such as the overlap with a GHZ state and the squeezing parameter, which are plotted as a function of $\gamma$ in Fig. 4.
@@
The perturbative regime for the XX and f-TAT models [see panels (b) and (c)] is supported by an appreciable overlap $\mathcal{F}_{\mathrm{GHZ}}$ with the GHZ state. However, unlike the infinite-range models, the overlap dips below a value of $1/2$ before transitioning to the squeezed-like regime. This shows that, while in the limit $\gamma\to 0$ the short-range interactions can generate quasi-collective entangled states (i.e., $\mathcal{F}_{\mathrm{GHZ}}>1/2$) Perlin et al. (2020), this eventually degrades as $\gamma$ is increased. Further to this, we highlight that the characteristic collectivity $\langle\hat{S}^{2}\rangle$ of the states generated in this regime is a substantial fraction of the maximum value $S(S+1)$. In contrast, the Ising model [see panel (a)] never features a meaningfully large overlap with the GHZ state, even for $\gamma\to 0$, although we note the value of the collectivity $\langle\hat{S}^{2}\rangle$ is not too different from the states prepared by XX and f-TAT.
@@
The squeezed-like regime between $\gamma^{(1)}<\gamma\leq\gamma^{(2)}$ is more complex for the short-ranged models. While the inverse squeezing parameter $(N\xi^{2}_{s})^{-1}$ takes non-zero values for $\gamma>\gamma^{(1)}$, reflecting the transition from the GHZ state to one with non-zero collective polarization of the qubits, in contrast to the results for the infinite-ranged models we observe that the squeezing parameter never breaches the SQL (i.e., $(N\xi^{2}_{s})^{-1}<1/N$). This reflects that while these states may exhibit anistropic fluctuations/correlations analogous to a spin-squeezed state(see averaged Husimi distributions in each panel), the finite-range interactions lead to states that are far from collective. Notably, the collectivity parameter $\langle\hat{S}^{2}\rangle$ is far smaller in the squeezed-like regime for these models than previous observations for the infinite-ranged models in Fig. 3. This complexity of the squeezed-like regime is further supported by computation of the CFI. For the Ising model, there is a small regime where the CFI breaches the SQL, but for the XX and f-TAT models the CFI is effectively always at or just worse than the SQL. Given the CFI that we compute is constrained to collective observables, this suggests that harnessing the QFI of states generated by finite-ranged interactions will require measurement of site-resolved observables or correlations.
Figure 5: Dephasing strengths $\gamma^{(j)}$ ($j=1,2$) as a function of $N$ for (a) OAT and TAT ($\alpha=0$), and (b) Ising, XX and f-TAT ($\alpha=6$). Solid and dashed lines are used to guide the eye and distinguish $\gamma^{(1)}$ and $\gamma^{(2)}$, respectively.
@@ III.3 Dependence on qubit number
We conduct additional calculations for system sizes $N=2,4,6,8$ and perform a similar analysis to that described in the previous sections. In particular, we still observe a general delineation of the optimal preparation strategy into the three regimes: perturbative, squeezed-like and uncorrelated. We extract the transition values $\gamma^{(1)}$ and $\gamma^{(2)}$ for each system size and plot the results in Fig. 5 (the solid and dashed lines are used to distinguish the markers for $\gamma^{(1)}$ and $\gamma^{(2)}$, respectively).
@@
Panel (a) of Fig. 5 shows the results for OAT and TAT. We observe that for both entangling Hamiltonians, the scope of the perturbative regime tends to decrease with system size, i.e., $\gamma^{(1)}$ trends to smaller values. This is unsurprising given that GHZ-like states are known to be highly fragile and difficult to dynamically prepare in the presence of decoherence, and this fragility worsens with system size Foss-Feig et al. (2013); Huelga et al. (1997). We note that while $\gamma^{(1)}$ appears to monotonically decrease with system size for OAT, TAT exhibits more sophisticated behaviour. Specifically, TAT depicts a slight increase in $\gamma^{(1)}$ from $N=2$ to $4$ before dropping away very quickly relative to OAT. We attribute this to the fact that TAT does not prepare explicit GHZ states (in fact the characteristic fidelity $\mathcal{F}_{\mathrm{GHZ}}$ deep in the perturbative regime changes noticeably as $N$ is varied) and we are probing particularly small system sizes. In future work, it would be interesting to exploit the permutational symmetry of the TAT dynamics using efficient numerical methods Baragiola et al. (2010) to study the non-unitary dynamics for much larger systems $N\gtrsim 100$, in an effort to better understand the complexity of the perturbative regime. In contrast to the shrinking perturbative regime, the value of $\gamma^{(2)}$ increases similarly with system size for OAT and TAT. Thus for large $N$ we expect that, for this VQC, preparing squeezed-like states will become the clear optimal strategy for sensing spin rotations in the presence of modest decoherence.
@@
Panel (b) of Fig. 5 shows analogous results for the Hamiltonians in Eqs. (2)-(4) with a choice of $\alpha=6$. Intriguingly, for the XX and f-TAT Hamiltonians we observe that $\gamma^{(1)}$ settles to a fixed value $\gamma^{(1)}\approx 0.46$ independent of system size for $N\geq 4$ (for the Ising Hamiltonian no perturbative regime exists so we do not plot $\gamma^{(1)}$). Note that this does not imply that finite-range interactions are more robust to decoherence, as the QFI of the states in this persistent perturbative regime can be much smaller than for infinite-range OAT and TAT (see Fig. 2). On the other hand, the increase of $\gamma^{(2)}$ with system size for each of the Ising, XX and f-TAT Hamiltonians closely follows the results for OAT and TAT.
IV Conclusion
With a focus on the preparation of states with optimal QFIs within a framework of noisy environments, our results confirm the existence of quantum states with QFIs above the SQL relevant to many fields, e.g., quantum metrology and quantum information science. We inform on the nature of the states generated with a set of physically relevant collective and finite-range Hamiltonians for a given decoherence $\gamma$ and classify them into three regimes closely associated with their QFIs and optimal preparation strategies. While we find that the generation of states with QFIs scaling near the HL appears to be an increasingly ambitious task with increasing system size (on account of rising sensitivities to deleterious noise), we emphasize that the generation of states with QFIs above the SQL scales favorably with $N$. Furthermore, our results indicate promise for the generation of states with QFIs practical toward quantum enhanced sensing and communication tasks with physical systems featuring limited interaction ranges.

### v2

@@ III.1.3 Additional diagnostic quantities
In addition to the fidelity to the GHZ state and the squeezing parameter, it is insightful to consider other quantities. First, we compute the collectivity of the prepared states (see Fig. 3), $\langle\hat{S}^{2}\rangle$, which should be compared against the maximal value of $S^{2}_{\mathrm{max}}=S(S+1)$ with $S=N/2$ for a collective state. Both OAT and TAT unitary dynamics preserve the collectivity since they feature infinite-range interactions, so that any deviation from the maximum value of $\langle\hat{S}^{2}\rangle$ should be attributable to the single-qubit dephasing noise. The collectivity is most useful when examined alongside a second quantity, the classical Fisher information (CFI), which can be obtained as
$F_{c}(\theta)=\lim_{\theta\to 0}\frac{4}{\theta^{2}}\sum_{m_{z}}\left(\sqrt{P(m_{z};\theta)}-\sqrt{P(m_{z};\theta=0)}\right)^{2},$ (10)
@@
where $P(m_{z})$ is the distribution function for the collective spin projection $m_{z}$ along $\hat{z}$ and $F_{c}$ is bounded from above by the QFI, $F_{c}\leq F_{Q}$. In the case of unitary OAT and TAT dynamics, the CFI always saturates the QFI when optimized over all possible collective measurements. However, we observe that $F_{c}$ is less than the QFI across parts of both the cat-like and squeezed-like regimes, particularly when the collectivity strays appreciably from $S^{2}_{\mathrm{max}}$. This suggests that more sophisticated measurements, such as local (site-resolved) spin observables, are necessary to exhaust the metrological potential of the states we preserve and, in particular, highlights that states in the “squeezed” regime—a name that is motivated by a simple analysis of collective spin fluctuations—require more nuanced treatment in general.
III.2 Finite-range interactions: Ising, XX, f-TAT
@@
Figure 4: Diagnostic quantities calculated for optimally prepared states as functions of relative dephasing strength $\gamma/\chi$ for (a)–(b) Ising, (c)–(d) XX, and (e)–(f) f-TAT Hamiltonians. All data is calculated for $N=8$ qubits. Panels (a), (c), and (e) are generated with $\alpha=3$ whereas panels (b), (d), and (f) use $\alpha=6$. Top subplots: Normalized classical Fisher information ($F_{c}/N^{2}$, squares), inverse Wineland squeezing parameter $(N\xi^{2}_{s})^{-1}$, tri-prongs), and the normalized QFI ($F_{Q}/N^{2}$, circles, diamonds ,and triangles, respectively). Lower subplots: Overlap of the optimal state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ with a GHZ state ($\mathcal{F}_{\mathrm{GHZ}}$, diamonds) and the normalized collectivity ($\langle\hat{S}^{2}\rangle/S^{2}_{\mathrm{max}}$, crosses). Vertical lines in all panels indicate $\gamma^{(1)}/\chi$ [absent for (a) and (b)] and $\gamma^{(2)}/\chi$. Below each set of panels we show Husimi phase space distributions for representative quantum states in each regime (see Appendix A for details of distribution). Bright regions indicate larger probability density. For $\gamma/\chi=0.01$ we show two perspectives of the same quantum state, rotated for indicated axes and angles.
@@
Equipped with an understanding of the results for the infinite-ranged OAT and TAT entangling Hamiltonians, we now generalize to the finite-ranged models given in Eqs. (2)–(4). For simplicity, wefocus on the specific coupling exponents $\alpha=3$ and $6$ in the following.
@@
In Figs. 2(b) and (c), we plot the maximal QFI and optimal parameters $\mathbf{x}_{\mathrm{opt}}$ for these coupling exponents as a function of dephasing strength $\gamma$ with $N=8$ qubits. Overall, we observe a similar delineation of the QFI and $\mathbf{x}_{\mathrm{opt}}$ into different regimes set by the dephasing strength, with some important caveats. First, for the XX and f-TAT models, we can define an analogous cat-like regime for $\gamma<\gamma^{(1)}$ (see vertical lines in panels), where the transition $\gamma^{(1)}/\chi$ is again defined by an abrupt drop in the interaction time characterized by $\theta_{I}$ and similarly sharp changes in the rotation angles/axes. Similarly, we define a squeezed-like regime for $\gamma^{(1)}\leq\gamma<\gamma^{(2)}$, with the upper bound on the dephasing strength $\gamma^{(2)}/\chi$ set via the uncorrelated regime that emerges when $\theta_{I}$ finally reaches zero. On the other hand, the Ising model presents distinct behavior. In particular, we observe an absence of the cat-like regime, as the VQC parameters change smoothly throughout a large squeezed-like regime $\gamma<\gamma^{(2)}$, until $\theta_{I}$ vanishes at the boundary of the uncorrelated regime that exists for $\gamma\geq\gamma^{(2)}$. We note that the extent of the cat-like (where applicable) and squeezed-like regimes is slightly reduced with $\alpha=3$ when compared with $\alpha=6$.
@@
While our preliminary identification of the cat-like, squeezed-like, and uncorrelated regimes is supported by an analysis of the optimal parameters $\mathbf{x}_{\mathrm{opt}}$, which show some qualitative similarities to OAT (Ising and XX) and TAT (f-TAT), we provide stronger evidence through quantities such as the overlap with a GHZ state and the squeezing parameter, which are plotted as a function of $\gamma/\chi$ in Fig. 4.
@@
The cat-like regime for the XX and f-TAT models [see panels (c)-(f) of Fig. 4] is supported by an appreciable overlap with the GHZ state Perlin et al. (2020). In particular, for $\gamma/\chi\to 0$ both XX and f-TAT feature $\mathcal{F}_{\mathrm{GHZ}}>1/2$, indicating the generation of a quasi-collective state that shows signatures consistent with a GHZ-like state. Unlike the infinite-range TAT results, however, the overlap for f-TAT dips below the value of $1/2$ before the transition to the squeezed-like regime. We also observe that the maximum value of $\mathcal{F}_{\mathrm{GHZ}}$ and indeed the QFI $F_{Q}/N^{2}$ attained for the XX and f-TAT Hamiltonians are surprisingly greater in the case $\alpha=6$ rather than the longer-range case with $\alpha=3$. This distinction is discussed in further detail momentarily. In addition to the GHZ fidelity, we highlight that the characteristic collectivity $\langle\hat{S}^{2}\rangle$ of the states generated in the cat-like regime remains a substantial fraction of the maximum value $S^{2}_{\mathrm{max}}$. In contrast, the Ising model [see panels (a) and (b) of Fig. 4] never features a meaningfully large overlap with the GHZ state, even for $\gamma/\chi\to 0$, although we note that the value of the collectivity $\langle\hat{S}^{2}\rangle$ remains appreciable, much like the states prepared by XX and f-TAT. Furthermore, unlike the XX and f-TAT examples, for the Ising Hamiltonian the attainable QFI in this regime is marginally reduced for $\alpha=6$ relative to that of $\alpha=3$.
@@
The squeezed-like regime between $\gamma^{(1)}\leq\gamma<\gamma^{(2)}$ is also significantly more complex for the finite-ranged models. While the inverse squeezing parameter $(N\xi^{2}_{s})^{-1}$ continues to take on non-zero values for $\gamma\geq\gamma^{(1)}$, reflecting the transition from the GHZ state to one with non-zero collective polarization of the qubits, in contrast to the results for the infinite-ranged models we observe that the squeezing parameter only marginally breaches the SQL (i.e., $(N\xi^{2}_{s})^{-1}<1/N$) for limited values of $\gamma$. This reflects that while these states may exhibit anisotropic fluctuations/correlations analogous to a spin-squeezed state (see averaged Husimi distributions shown beneath each panel in Fig. 4), the finite-range interactions lead to states that are far from collective. Notably, the collectivity parameter $\langle\hat{S}^{2}\rangle$ is far smaller in the squeezed-like regime for these models than previous observations for the infinite-ranged models in Fig. 3.
The complexity of the squeezed-like regime is further supported by computing the CFI. For the Ising model, the CFI lies above the SQL for $\gamma\ll\gamma^{(2)}$ but quickly deviates below the QFI as $\gamma$ increases. On the other hand, for the XX and f-TAT models the CFI shows only a handful of points in the squeezed-like regime where the SQL is breached and does not approach the QFI until the point at which quantum enhancement is essentially lost. Given that the CFI we compute is constrained to collective observables, this suggests that, in general, harnessing the QFI of states generated by finite-ranged interactions will require measurement of site-resolved observables or correlations.
Figure 5: Optimal QFI (top subplots), diagnostic quantities (middle subplots), and interaction parameters $\theta_{\mathrm{I}}$ (lower subplots) as functions of interaction range $\alpha$ for Ising (blue circle), XX (green diamond), and f-TAT models (red triangle). All points are computed with $N=8$ for choices of (a) $\gamma/\chi=0.01$ and (b) $\gamma/\chi=1.0$. Note that we present the QFI and inverse squeezing quantity on a different vertical scale in panel (b).
We carry out a more systematic investigation of the influence of the interaction range in Fig. 5, wherein we plot the optimal QFI for each choice of entangling Hamiltonian as a function of $\alpha$. Panels (a) and (b) present results for fixed values of the decoherence rate $\gamma/\chi=0.01$ and $\gamma/\chi=1.0$, respectively. To further characterize the generated states, we plot the corresponding VQC parameter $\theta_{I}$ and (a) the fidelity with a GHZ state ($\mathcal{F}_{\mathrm{GHZ}}$) or (b) the inverse squeezing parameter ($(N\xi^{2}_{s})^{-1}$). All data uses $N=8$ qubits.
In the case of weak decoherence [Fig. 5(a), $\gamma/\chi=0.01$], the QFI behaves in a markedly distinct manner for each of the models. In the case of the Ising Hamiltonian, the QFI begins near the HL ($F_{Q}/N^{2}\approx 1$) for $\alpha=0$ and quickly decreases as $\alpha$ increases (i.e., the interaction range shrinks), approaching a value of $F_{Q}/N^{2}\approx 0.27$ (still just above the SQL, $F_{Q}/N^{2}=1/8$) for $\alpha=6$. The fidelity with a GHZ state (middle panel) similarly drops as the interaction is decreased, which explains the reduction of the QFI relative to the HL. Underpinning both observations, the VQC parameter $\theta_{I}$ abruptly drops away from values associated with the generation of GHZ-like states ($\theta_{I}/\pi\sim 0.5$) toward faster evolution that indicates the optimal VQC is preparing squeezed-like states.
On the other hand, while for the XX and f-TAT cases the QFI again starts near the HL for $\alpha=0$, it does not monotonically decrease. Instead, it features a dip for intermediate values of $1\lesssim\alpha\lesssim 3$ before growing again to become near the HL for the larger values of $\alpha$ we probe (reaching $F_{Q}/N^{2}\approx 0.83$ for $\alpha=6$). Despite the qualitative similarity, the behavior of $\theta_{I}$ provides insight that distinguishes the two Hamiltonians. For the XX Hamiltonian, we observe that $\theta_{I}$ is relatively unchanged as $\alpha$ is varied, suggesting that the existence of the cat-like regime is robust regardless of interaction range Perlin et al. (2020). This assertion is supported the fact that the overlap with the GHZ state remains appreciable throughout (i.e., $\mathcal{F}_{\mathrm{GHZ}}$ is above or close to $1/2$ throughout). By contrast, for the f-TAT Hamiltonian the more substantial dip in the QFI is accompanied by an abrupt drop in the interaction duration $\theta_{I}$ to smaller values consistent with the generation of squeezed-like states, before jumping back to larger values consistent with the generation of cat-like states. Again, this behavior is supported by a computation of $\mathcal{F}_{\mathrm{GHZ}}$, which similarly falls below $1/2$ in the region of $\alpha$ where smaller $\theta_{I}$ is obtained before rising back above $1/2$ when $\theta_{I}$ increases. The behavior of the QFI at large $\alpha$ for both the XX and f-TAT Hamiltonians suggests that the generation of quasi-collective, cat-like states with large QFI is possible with short-range interactions (at least in small systems) and merits further investigation.
The case of more moderate decoherence [Fig. 5(b), $\gamma/\chi=1.0$] presents a simpler story, with the QFI starting from smaller values at $\alpha=0$ and monotonically decreasing as $\alpha$ is increased. In parallel, the values of the associated optimal VQC parameter $\theta_{I}$ (and the inverse squeezing parameter $(N\xi^{2}_{s})^{-1}$) are consistent with those previously observed in the squeezed-like regime, suggesting it remains robust regardless of $\alpha$.
@@ III.3 Dependence on qubit number
We conduct additional calculations for system sizes $N\in\{2,4,6,8\}$ and perform a similar analysis to that described in the preceding sections. In particular, we still observe a general delineation of the optimal preparation strategy into the three regimes: cat-like, squeezed-like,, and uncorrelated. We extract the transition values $\gamma^{(1)}/\chi$ and $\gamma^{(2)}/\chi$ for each system size and plot the results in Fig. 6 (the solid and dashed lines are used to distinguish the markers for $\gamma^{(1)}/\chi$ and $\gamma^{(2)}/\chi$, respectively).
Figure 6: Dephasing strengths $\gamma^{(j)}$ ($j=1,2$) as a function of $N$ for (a) OAT and TAT with $\alpha=0$, and Ising, XX, and f-TAT with (b) $\alpha=3$ and (c) with $\alpha=6$. Solid and dashed lines depict $\gamma^{(1)}/\chi$ and $\gamma^{(2)}/\chi$, respectively.
@@
Panel (a) of Fig. 6 shows the results for OAT and TAT. We observe that for both entangling Hamiltonians, the scope of the cat-like regime tends to decrease with system size; in other words, $\gamma^{(1)}/\chi$ trends to smaller values. This is unsurprising given that GHZ-like states are known to be highly fragile and difficult to dynamically prepare in the presence of decoherence, and this fragility worsens with system size Foss-Feig et al. (2013); Huelga et al. (1997). We note that while $\gamma^{(1)}/\chi$ appears to monotonically decrease with system size for OAT, TAT exhibits more sophisticated behavior. Specifically, TAT depicts a slight increase in $\gamma^{(1)}/\chi$ from $N=2$ to $4$ before dropping away quickly relative to OAT. We attribute this to the fact that TAT does not prepare explicit GHZ states (in fact the characteristic fidelity $\mathcal{F}_{\mathrm{GHZ}}$ deep in the cat-like regime changes noticeably as $N$ is varied) and we are probing particularly small system sizes. In future work, it would be interesting to exploit the permutational symmetry of the TAT dynamics using efficient numerical methods Baragiola et al. (2010) to study the non-unitary dynamics for much larger systems $N\gtrsim 100$, in an effort to better understand the complexity of the cat-like regime. In contrast to the shrinking cat-like regime, the value of $\gamma^{(2)}/\chi$ increases similarly with system size for OAT and TAT. Thus, for large $N$ we expect that, for this VQC, preparing squeezed-like states will become the clear optimal strategy for sensing spin rotations in the presence of modest decoherence.
@@
Panels (b) and (c) of Fig. 6 depict analogous results for the Hamiltonians in Eqs. (2)–(4) with a choice of $\alpha=3$ and $\alpha=6$, respectively. Intriguingly, in panel (c) for the XX and f-TAT Hamiltonians we observe that $\gamma^{(1)}/\chi$ settles to a fixed value $\gamma^{(1)}/\chi\approx 0.46$ independent of system size for $N\geq 4$ (for the Ising Hamiltonian no cat-like regime exists so we do not plot $\gamma^{(1)}/\chi$). Note that this does not imply that finite-range interactions are more robust to decoherence, since the QFI of the states in this persistent cat-like regime can be much smaller than for infinite-range OAT and TAT (see Fig. 2). On the other hand, the increase of $\gamma^{(2)}/\chi$ with system size for each of the Ising, XX, and f-TAT Hamiltonians closely follows the results for OAT and TAT.
IV Optimal state preparation with general decoherence
In our investigation so far, we have focused on a relatively specific scenario of isotropic dephasing. Here we broaden the scope of our work and present general results for (i) imbalanced dephasing along each axis and (ii) an alternative scenario where the decoherence is due to a combination of spontaneous emission and dephasing along the $z$-axis. While the presence of the cat-like, squeezed-like, and uncorrelated regimes remains, we find that the quantitative extent of these regimes is sensitively dependent on the specific combination of the entangling Hamiltonian and dominant decoherence mechanism.
We first present results for the case of $\alpha=0$ (i.e., OAT and TAT Hamiltonians with infinite-range interactions) with $N=6$ qubits in Fig. 7. Panels (a) and (b) present the optimal QFI ($F_{Q}/N^{2}$) for OAT and TAT, respectively, as a function of both $z$-dephasing with strength $\gamma_{z}/\chi$ and $y$-dephasing with strength $\gamma_{y}/\chi$ (we ignore $x$-dephasing by setting $\gamma_{x}=0$). [^1] Note that we do not present similar depictions for $x$- versus $z$-dephasing and $x$- versus $y$-dephasing as both $x$- and $y$-dephasing render identical results. A similar statement can be made for the interchanging of spontaneous emission/absorption. To guide the eye, we indicate the boundaries between the cat-like and squeezed-like regimes with magenta lines and similarly between the squeezed-like and uncorrelated regimes with green lines. The boundaries are defined analogously to $\gamma^{(1)}$ and $\gamma^{(2)}$ indicated previously in the results of Sec. III.
Figure 7: Optimal QFI as a function of relative strength of various types of decoherence for OAT in panels (a) and (b) TAT in panels (b) and (d). All data is for $N=6$ and $\alpha=0$. The top row scans over $z$-dephasing ($\gamma_{z}/\chi$) versus $y$-dephasing ($\gamma_{y}/\chi$) whereas the bottom row scans over $z$-dephasing versus spontaneous emission ($\gamma_{-}/\chi$). The boundary between the cat-like/squeezed-like regimes and squeezed-like/uncorrelated regimes is delineated by magenta and green lines, respectively.
Starting with OAT in Fig. 7(a), we observe that the extent of the cat-like regime is nearly symmetric, with only a slightly greater sensitivity to $y$-dephasing than $z$-dephasing. In contrast, we observe that the squeezed-like regime is very asymmetric and is clearly less sensitive to $y$-dephasing. Specifically, for small $\gamma_{y}/\chi\ll 1$ we find the SQL is reached for $\gamma_{z}/\chi\sim 10^{1}$, whereas in the opposing case with small $\gamma_{z}/\chi\ll 1$ the squeezed-like regime extends far beyond $\gamma_{y}/\chi\gg 10^{1}$ [i.e., extends outside the range shown in panel (a)]. In the case of TAT [see 7(b)], while we find that the cat-like regime is marginally more sensitive to $z$-dephasing, the more striking observation is the reduced extent of this regime relative to OAT. On the other hand, the squeezed-like regime is much larger than that observed for OAT and notably features larger QFI at comparable values of decoherence. Both observations are consistent with earlier data shown in Fig. 3 for isotropic dephasing.
We additionally present results for OAT and TAT subject to spontaneous emission and $z$-dephasing in Fig. 7(c) and (d), where we show the QFI as a function of the respective decoherence rates $\gamma_{-}/\chi$ and $\gamma_{z}/\chi$. In the case of OAT, while the cat-like regime features a slightly better robustness to $z$-dephasing as opposed to spontaneous emission, the squeezed regime clearly showcases the reverse. The onset of the SQL occurs near $\gamma_{z}/\chi\approx 10$ when $\gamma_{-}/\chi$ is taken to be vanishingly small; in contrast, when $\gamma_{z}/\chi$ is negligible, the squeezed regime continues to persist for large spontaneous emission $\gamma_{-}/\chi>10^{1.5}$ (the crossover the to SQL is not shown). For TAT [see 7(d)], the extent of the cat-like regime is contracted relative to OAT with the boundary values of $\gamma_{-}/\chi$ and $\gamma_{z}/\chi$ (magenta line) effectively reduced by an order of magnitude compared with panel (c). However, the squeezed-like regime again encompasses a considerably larger range of decoherence rates relative to OAT. Specifically, the onset of the uncorrelated regime occurs at large values of $\gamma_{z}/\chi,\gamma_{-}/\chi>10^{1.5}$, which are outside the range of parameters displayed in panel (d). In addition, by inspecting the value of the QFI within the squeezed-like regime, we identify a relative robustness to $z$-dephasing: The onset of red hues (lower QFI) occurs in the limiting case of dominant dephasing at ($\gamma_{-}/\chi,\gamma_{z}/\chi)\approx(10^{-2},10^{1})$ as opposed $(\gamma_{-}/\chi,\gamma_{z}/\chi)\approx(10^{0},10^{-2})$ in the case of dominant spontaneous emission.
Figure 8: Optimal QFI as a function of relative strength of various types of decoherence for Ising [panels (a) and (e)], XX [panels (b) and (e)], and f-TAT [panels (c) and (f)]. All data is for $N=6$ and $\alpha=6$. The top row scans over $z$-dephasing ($\gamma_{z}/\chi$) versus $y$-dephasing ($\gamma_{y}/\chi$), and the bottom row scans $z$-dephasing versus spontaneous emission ($\gamma_{-}/\chi$). The boundary between the cat-like/squeezed-like regimes and squeezed-like/uncorrelated regimes are delineated by magenta and green lines, respectively.
Figure 8 presents analogous results for Ising, XX, and f-TAT models with $\alpha=6$ and $N=6$ qubits. Panels (a)–(c) display the optimized QFI as a function of dephasing rates $\gamma_{y}/\chi$, and $\gamma_{z}/\chi$ ($\gamma_{x}$ is again set to zero). For the XX and f-TAT models, the cat regime is of similar size and approximately symmetric with respect to the dephasing axis. On the other hand, the Ising model does not feature any cat regime, consistent with our prior discussion in Sec. III. Similar to the $\alpha=0$ results for OAT, the squeezed-like regime of the Ising and XX models extends to large values of $\gamma_{y}/\chi$ falling outside the plot range and is much more robust to $y$-dephasing. Equivalently, the squeezed-like regime for the f-TAT model qualitatively mirrors the $\alpha=0$ results for TAT, being far more robust than the Ising and XX models.
Panels (d)–(f) of Fig. 8 illustrate the behavior of the optimal QFI as a function of $z$-dephasing and spontaneous emission rates $\gamma_{z}/\chi$ and $\gamma_{-}/\chi$. Overall, the relative size and existence of the cat-like and squeezed-like regimes follow similar trends to the data of panels (a)–(c). The most significant difference in the case of imbalanced dephasing is the relatively quicker dropoff of the QFI for the XX model when spontaneous emission is dominant, while the converse is true for f-TAT.
V Conclusion
Our work provides insight into optimal state preparation strategies to maximize the QFI for metrology tasks while fully accounting for limitations imposed by decoherence. In particular, we identify a clear delineation of the strategies in terms of the types of states they generate, regardless of the entangling interactions used as part of the VQC. Moreover, while we find that the generation of states with QFI scaling near the HL appears to become increasingly ambitious for larger system sizes, as a result of the increased fragility of the preparation schemes to decoherence, we emphasize that the generation of states with a QFI above the SQL can scale favorably with $N$. Furthermore, our results suggest that finite-range interactions may be sufficient to generate states with appreciable QFI.